The rabbit population in a Victorian town was estimated to be 320,000 at the start of 2012. Scientists believe that this will increase by two percent each year. What will the rabbit population be at the end of 2015?



1
Expert's answer
2021-02-24T07:08:07-0500
"N(n)=N_0(1+r)^n"

"N_0=320000" is the rabbit population in a Victorian town at the start of 2012.

"r=0.02" is the growth rate

"n=2016-2012=4" is the number of yers from the the start of 2012 to the end of 2015.


"N(3)=320000(1+0.02)^4=346378"

The rabbit population in a Victorian town will be 346378 at the end of 2015.